OpenStudy (anonymous):

Find the diameter of a cone that has a volume of 104.67 cubic inches and a height of 4 inches. Use 3.14 for pi. 4 inches 6 inches 8 inches 10 inches

OpenStudy (anonymous):

the first thing is the formula, what is the formula?

OpenStudy (anonymous):

Im not sure

OpenStudy (anonymous):

let that equal your volume i.e. \[pixr^2x (h/3) = 104.67 cm^3\]

OpenStudy (anonymous):

Now, you are given the height = 4 inches so, \[\pi x r^2 x (4/3) = 104.67 cm^3\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

now, lets isolate the r^2

OpenStudy (anonymous):

so \[\pi x (4/3) = \]

OpenStudy (anonymous):

roughly 4.18879

OpenStudy (anonymous):

so now you have 4.18879 r^2 = 104.67 cm^3

OpenStudy (anonymous):

dividing across, we get \[104.67cm^3/4.18879 = 24.988 cm^3\]

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

r^2 = 24.988cm^3

OpenStudy (anonymous):

Sorry im so confused with this question

OpenStudy (anonymous):

\[r= \sqrt{24.988}\]

OpenStudy (anonymous):

r= 4.99

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

So its ten

OpenStudy (anonymous):

try it now: \[V= \pi x r^2 x (h/3)\]

OpenStudy (anonymous):

\[V= \pi x (5)^2 x (4/3)\]

OpenStudy (anonymous):

Now since r= 5, and d= 2r, then the diameter must be 10inches

OpenStudy (anonymous):

Yes i get it now thank you so much

OpenStudy (anonymous):

You're very welcome!